# Probing vector chirality in the early Universe

### v1

@@ Abstract
We explore the potential of detecting parity violation in primordial vector fossils using late-time galaxy spins. Utilizing $N$-body simulations, we use halo spins as a reliable proxy for galaxy spins to investigate how effectively such primordial vectorial parity asymmetry remains in galaxy spins at low redshifts. We develop a novel approach to generate initial conditions with substantial parity asymmetry, while maintaining the initial matter power spectrum unchanged. From the parity broken initial condition and halos evolved from it, we construct the initial spin and halo spin fields, respectively. Focusing on the helicity of these vector fields, we detect substantial asymmetry in the initial spin field as a consequence of parity violation in the primordial vector fossil. In addition, we discover that over $50\%$ of the primordial asymmetry in the initial spin field remains in the late-time halo spin field on a range of scales. Given the tight correlation between halo spins and observable galaxy spins, we expect to detect the current amplitude of vectorial parity asymmetry potentially up to $16\sigma$-level in observation, when utilizing galaxy samples from DESI BGS. Our findings demonstrate that the primordial imprints of vectorial parity violation persist through non-linear gravitational evolution, highlighting the reliability of galaxy spin as a sensitive probe for testing the vectorial parity-invariance in the early Universe.
@@
Recently, unexpected $3-7\sigma$ detections of parity asymmetry in the observed galaxy distribution were reported [1, 2] focusing on the “helicity” (or handedness) of galaxy quartets in 3-dimension. Although particular care must be taken for accurately modeling systematic uncertainties to conclude such signal to be of primordial origin [2, 3, 4, 5, 6] this motivates further exploration for various types of potential parity violation (PV) in the early Universe and their signatures.
@@
Primary focus has been on the theoretical search for tensorial type, where some inflationary models can produce unequal primordial gravity waves (pGWs) with left- and right-helicities [7, 8, 9, 10, 11, 12]. Thus, detecting either the helical asymmetry between two chiral-pGWs or their imprints on CMB polarization [13, 14, 15, 16] and LSS [17, 18, 19, 20, 21] could confirm the primordial tensorial-PV. Such helical asymmetry imprints would be preferably more detectable in 3-dimensional probes [18] since the helicity information is largely lost in 2-dimension, for example, leaving the detection in the polarized CMB [15, 22, 23, 24] more challenging [c.f. see 25, 26, 27, hinting at the non-vanishing EB-correlation of non-primordial origin, e.g., cosmic birefringence].
@@
Another interesting, but largely unexplored, class of parity violation is vectorial type, possibly generating helically asymmetric vector fossils [17] that may leave imprints in initial vector quantities. Recently, a 3-dimensional estimator “initial spin” (rigorously defined in Eq. 9) – a vector representing the spinning of a mass element torqued by the initial tidal field [28, 29, 30] – was proposed [31, 32] and examined [33] to search for vectorial helical asymmetry. This proposal relies on the two key properties of the initial spin field: 1) a vector field that only consists of left- and right-helical modes, and 2) its strong alignments with “galaxy spin” and “halo spin”, or the angular momentum direction of a galaxy [32, 34] and a dark matter halo [35, 36, 37, 31] at late-time. The latter suggests that the characteristics of the initial spin field will remain in the galaxy spin field. Therefore, detecting helical asymmetry in a galaxy spin field would indicate that the initial spin field is helically asymmetric, evidencing vectorial-PV in the early Universe.
@@
On the observational side, there has not yet been compelling evidence that the galaxy spin field is helically asymmetric; e.g., only discovered were its mildly uneven correlations to the left- and right-helical components of the reconstructed initial spin field [33]. However, it is not straightforward to conclude the absence of primordial vectorial-PV from such a weak-to-none helical asymmetry detection. This is because the primordial helical asymmetry could have been non-negligibly erased during galaxy formation and evolution, leaving only an insignificant signal, or near parity symmetry, in the galaxy spin. This necessitates a further investigation into how effectively the primordial helical asymmetry in the initial spin is carried onto the late-time galaxy spin.
@@
In this Letter, we investigate the detectability of parity violation in a primordial vector fossil, namely “helical” (i.e., rotational) displacement vector field, using galaxy spins. The helical displacement vector field is envisioned to be sourced in the very early universe, e.g., during inflation, and remains as fossils [17]. We emphasize that this vector fossil is not the initial vector mode perturbations, which decay over time. Utilizing numerical N-body simulations, we use halo spins as a reliable proxy for galaxy spins, leveraging the tight correlation between those two spins [38, 39, 37, 34]. In the spirit of effective field theory, without prior assumptions on microscopic physics, we develop a novel way of generating the initial condition with a helically asymmetric initial spin field. We show that broken parity symmetry in the helical displacement fossil efficiently produces strong primordial helical asymmetry without altering the initial matter power spectrum. By analyzing the spins of halos evolved from such initial conditions we demonstrate that such primordial helical asymmetry significantly survives in the final halo spin field.
I Simulation and Initial condition
@@
We construct 500 pairs of N-body simulations – ones with parity symmetric initial condition and the others with parity asymmetric initial condition. The simulations adopt the modified version of the code CUBE [40] to evolve $256^{3}$ dark matter particles within a periodic cubic box of size $L=100~{}h^{-1}{\rm Mpc}$. For numerical convenience, the initial redshift is set $z_{\rm init}=30$. We generate the initial density fluctuation of the standard parity symmetric universe using Zel’dovich approximation [41] where the displacement vector field, $\bm{\Psi}_{\rm sy}$, is curl-free (i.e., irrotational) and its Fourier space representation follows
$\bm{\Psi}_{\rm sy}=\bm{\Psi}_{\rm ZA}=-i\bm{k}\frac{\Phi(\bm{k})}{4\pi Ga^{2}\bar{\rho}}\ ,$ (1)
@@
with the gravitational potential $\Phi$, scale factor $a$, gravitational constant $G$, and mean cosmic density $\bar{\rho}$. On the other hand, the parity asymmetric initial condition is constructed using the modified displacement field, $\bm{\Psi}_{\rm asy}$, including an extra helical displacement, $\bm{\Psi}_{\rm hel}$,
$\bm{\Psi}_{\rm asy}=\bm{\Psi}_{\rm ZA}+\bm{\Psi}_{\rm hel}.$ (2)
@@
$\bm{\Psi}_{\rm hel}$ $=\bm{\Psi}_{\mathcal{R}}+\bm{\Psi}_{\mathcal{L}}=\sum_{\mathcal{H}=\mathcal{R},\mathcal{L}}\Phi_{\mathcal{H}}(\bm{k})\hat{\bm{e}}_{\mathcal{H}}(\bm{k}),$ (3)
where the respective helical displacement field, $\bm{\Psi}_{\mathcal{R}/\mathcal{L}}$, can be constructed as the product of a pseudo-scalar field, $\Phi_{\mathcal{R}/\mathcal{L}}$, and an eigenvector of the curl operator in Fourier space,
$\hat{\bm{e}}_{\mathcal{R}/\mathcal{L}}(\bm{k})=\frac{1}{\sqrt{2}k}\frac{1}{\sqrt{k_{y}^{2}+k_{z}^{2}}}\begin{pmatrix}k_{y}^{2}+k_{z}^{2}\\ -k_{x}k_{y}\pm ikk_{z}\\ -k_{x}k_{z}\mp ikk_{y}\end{pmatrix}\ ,$ (4)
with $k=|\bm{k}|=\sqrt{k_{x}^{2}+k_{y}^{2}+k_{z}^{2}}$. The parity symmetry in the primordial vector fossil is broken if the modified displacement field includes two helical components unequally. To simplify our analysis, we assume $\bm{\Psi}_{\rm hel}$ to be purely right-handed, e.g., $\bm{\Psi}_{\rm asy}=\bm{\Psi}_{\rm ZA}+\bm{\Psi}_{\mathcal{R}}$. Note that both parity symmetric and asymmetric simulations should have identical initial density perturbation in Lagrangian space. This is because the initial density perturbation is solely generated by the same $\bm{\Psi}_{\rm ZA}$ [42], but not by $\bm{\Psi}_{\rm hel}$, as the latter is divergence-free. However, the helical displacement only re-maps the density perturbation at one (Eulerian) position to a different position and plays a pivotal role in breaking parity symmetry in the initial spin vector field, as we will show later.
We now describe the generation of $\Phi$ and $\Phi_{\mathcal{R}}$. We calculate $\Phi$ from the linear matter power spectrum $P(k)$,
$\langle\delta(\bm{k})\delta^{*}(\bm{k}^{\prime})\rangle=(2\pi)^{3}\delta_{D}(\bm{k}-\bm{k}^{\prime})P(k),$ (5)
via the Poisson equation $-k^{2}\Phi=4\pi Ga^{2}\bar{\rho}\delta$, where $\delta$ is the linear density fluctuation. The normalization amplitude is then defined
[…]
and set to $\sigma(z_{\rm init})=0.20$. Note that this corresponds to familiar $\sigma_{8}(z=0)=0.81$, consistent with CMB measurements [43]. Similarly, the pseudo-scalar field $\Phi_{\mathcal{R}}$ is obtained following
$\langle k^{2}\Phi_{\mathcal{R}}(\bm{k})k^{\prime 2}\Phi_{\mathcal{R}}^{*}(\bm{k}^{\prime})\rangle=(2\pi)^{3}\delta_{D}(\bm{k}-\bm{k}^{\prime})P_{\mathcal{R}}(k),$ (7)
where we assume $P_{\mathcal{R}}(k)$ adopts high-$k$ suppression on $P(k)$ as
@@
We set $R_{\rm h}=0.6~{}h^{-1}{\rm Mpc}$ in correspondence to the Gaussian radius of a Lagrangian protohalo of $10^{12}~{}h^{-1}{\rm M}_{\odot}$. The normalization is set to $\sigma_{\mathcal{R}}(z_{\rm init})=1.45$ to produce large enough parity asymmetry while maintaining the initial matter power spectrum unchanged. Qualitatively, $\sigma_{\mathcal{R}}$ determines the degree of primordial parity asymmetry, in analogy to non-gaussianity parameter, $f_{\rm NL}$, describing the departure from Gaussianity [44]. A universe without vectorial-PV corresponds to the case with $\sigma_{\mathcal{R}}=0$.
Independent random number sets are used for generating $\Phi$ and $\Phi_{\mathcal{R}}$ to make them uncorrelated. We identify halos using the Friends-of-Friends algorithm and utilize only halos more massive than $10^{12}~{}h^{-1}{\rm M}_{\odot}$. The total number of such halos in the entire 500 simulations of each kind is $N_{\rm halo}\simeq 10^{6}$.
II Initial density field
Figure 1: (Eulerian) Initial matter density fields at $z_{\rm init}=30$ in the parity symmetric (left) and asymmetric (right) simulations sliced in a $0.4~{}h^{-1}{\rm Mpc}$-thickness slab. Density fields are calculated using the could-in-cell (CIC) assignment scheme and smoothed with a Gaussian kernel over $R_{\rm h}=0.6h^{-1}{\rm Mpc}$. Note that they look identical in Lagrangian space.
We begin by visually comparing the initial density fields of the parity symmetric and asymmetric simulations in Fig. 1. It can be seen that the helical displacement field $\Psi_{\mathcal{R}}$ contributes to distorting the standard density field similarly to the gravitationally lensed map of the CMB temperature and polarization [45, 46, 47]. For example, the density structures in the parity asymmetric realization tend to be stretched and/or warped from their counterparts at similar locations in the standard parity symmetric realization. However, unlike CMB maps exhibiting concentric distortion around over/under-densities, the density field deformation due to the helical displacement occurs incoherently.
III Initial and halo spin fields
We now examine the parity symmetry of the initial conditions by measuring the helical asymmetry of the initial spin vector field. The initial spin field, $\bm{J}_{\rm init}$, is calculated on $256^{3}$ grids as a unit vector field constructed from the Hessians of the smoothed initial density, $\rho_{\rm init}^{s}$, and potential field, $\phi_{\rm init}^{s}$, [35, 48, 31, 32, 33],
@@ III Initial and halo spin fields
with the Levi-Civita symbol $\epsilon_{\alpha\beta\gamma}$. The optimal smoothing scale of the Gaussian kernel is set to $R_{\rm s}=2R_{\rm G}$, allowing us to achieve the maximum significance for detecting the primordial vectorial-PV. Note that the initial spin vector field is divergence-free, or purely helical, by construction. On the other hand, the spin vector of a dark matter halo is defined as the unit angular momentum vector of a halo at $z=0$,
$\bm{j}_{\rm halo}\propto\sum_{i}\bm{r}_{p}^{i}\times\bm{v}_{p}^{i},$ (10)
where $\bm{r}_{p}^{i}$ and $\bm{v}_{p}^{i}$ are the relative position and velocity vectors of the $i^{\rm th}$-member particle measured from its halo center. A halo spin field is then constructed by assigning these halo spin vectors to their protohalo centers, $\bm{x}_{\rm ph}$, at $z=z_{\rm init}$,
$\bm{J}_{\rm halo}=\begin{cases}\bm{j}_{\rm halo}&\text{if $\bm{x}=\bm{x}_{\rm ph}^{i}$ for $i^{th}$ halo}\\ \bm{0}&\text{otherwise}.\end{cases}$ (11)
Here, $\bm{x}_{\rm ph}$ is defined as the mass center of halo member particles at their initial positions. We also construct the subsampled initial spin field only using the initial spin vectors at the same protohalo positions $\bm{x}=\bm{x}_{\rm ph}^{i}$, similarly to the halo spin field as in Eq. 11.
IV Measuring Helical asymmetry
We examine the parity symmetry of the initial, subsampled initial, and halo spin fields by measuring the helical asymmetry between their $\mathcal{R}$- and $\mathcal{L}$-handed components. We decompose each of the constructed spin fields into its $\mathcal{R}$- and $\mathcal{L}$-helical fields so that $\bm{J}=\bm{J}_{\mathcal{R}}+\bm{J}_{\mathcal{L}}$. The $\mathcal{R}$$\mathcal{L}$-helical spin field can be computed via
$\bm{J}_{\mathcal{R/L}}=(J_{\mathcal{R/L}})_{\alpha}=\sum_{\beta}\mathbb{P}^{\mathcal{R/L}}_{\alpha\beta}({\bm{k}})J_{\beta}({\bm{k}})$ (12)
@@ IV Measuring Helical asymmetry
The helical asymmetry, $\chi$, of a spin field can be then quantified at each wavenumber $k$ following
@@
where $P_{{\mathcal{RR}/\mathcal{LL}}}(k)$ represents the power spectrum of the $\mathcal{R}$$\mathcal{L}$-helical spin field defined as
@@
A spin vector field with a perfect helical symmetry satisfies $\chi=0$, whereas $|\chi|$ equals unity when the spin field is maximally asymmetric, or purely $\mathcal{R}$$\mathcal{L}$-helical.
Figure 2: Helical asymmetry spectrum (upper) of initial spin field and the significance of the departure (lower) from the perfect symmetry in the parity symmetric and asymmetric simulations. The vertical line marks a wavenumber corresponding to twice the Gaussian radius of a $10^{12}~{}h^{-1}{\rm M}_{\odot}$ protohalo, $k=1/(2R_{\rm h})$, whereas the horizontal line indicates $\chi/\sigma=50$. Shaded regions represent the standard error of the mean calculated from the 500 realizations, or the total volume of $0.5(h^{-1}{\rm Gpc})^{3}$.
In Fig. 2, we show the helical asymmetry in the full initial spin field in the parity symmetric/asymmetric realizations. We focus only on scales larger than $k=1/(2R_{\rm h})$, below which the helical asymmetry feature in halo spins is likely to be convolved with the halo exclusion effect [49]. In the parity asymmetric realizations, the initial spin field exhibits scale-dependent helical asymmetry that deviates from $\chi=0$, maximally reaching $70\sigma$ detection. Overall, the initial spin field of the parity asymmetric realization is preferentially $\mathcal{R}$-handed, e.g., $\chi>0$. In particular, the helical asymmetry can be as large as $\chi_{\rm peak}\approx 0.18$ indicating that the $\mathcal{R}$-helical mode is stronger by $44\%$ than the $\mathcal{L}$-helical mode. Interestingly, the relative dominance between the two helical modes of the initial spin field is determined to be identical to that of the helical displacement field. On the other hand, as expected, we observe a nearly perfect helical symmetry (i.e., $\chi\approx 0$) on all scales investigated in the parity symmetric simulations.
@@
We now compare the late-time halo spin field with the subsampled initial spin field in Fig. 3. Note that we rescaled all helical asymmetry spectra by multiplying the same rescaling factor, which makes the subsampled and full initial spin fields for the parity asymmetric case have identical peak height, e.g., $\chi_{\rm peak}\simeq 0.18$. In the parity symmetric cases, the subsampled initial field shows larger fluctuations around $\chi=0$ than in the full initial spin field. Nonetheless, it is still consistent with the helical symmetry within $1.4\sigma$ significance. Similarly, we do not detect any significant deviation from the perfect symmetry in the halo spin field. The departure at maximum is only about $0.7\sigma$-level across the scales investigated. This indicates that the halo spin field remains parity symmetric if the initial spin field were parity symmetric.

### v2

@@ Abstract
We explore the potential of using late-time galaxy spins to test the parity symmetry of primordial vector fossils. Using $N$-body simulations, we analyze halo spins as a reliable proxy for galaxy spins to investigate the detectability of this effect. We develop a novel approach to generate initial conditions (ICs) that have substantial parity asymmetry but do not alter the initial matter power spectrum. We construct the initial spin fields from the parity broken ICs and halo spin fields using late-time halos evolved from such ICs. Focusing on the helicity of these vector fields, we detect substantial asymmetry in the initial spin field. In addition, we find that over $50\%$ of the initial spin field’s asymmetry remains in the late-time halo spin field on a range of scales. Based on mock galaxy spin fields derived from the halo spin fields, we forecast that a maximum detection at $13\sigma$ is possible with the final DESI BGS for the model considered in this analysis. Our findings demonstrate that primordial vectorial parity violation survives nonlinear gravitational evolution, and thus, can be effectively probed with galaxy spins at late times.
@@
Recently, unexpected $3-7\sigma$ signals of parity asymmetry were detected [1, 2] in observation using parity-odd basis [3, 4], with a focus on the “helicity” (or handedness) of galaxy quartets in 3 dimensions. Although particular care must be taken for accurately modeling systematic uncertainties to conclude such a signal to be of primordial origin [5, 2, 6, 7, 8, 9], this motivates further exploration for various types of potential parity violation (PV) in the early Universe and their signatures.
@@
The primary focus has been on the theoretical search for tensorial type, where some inflationary models can produce unequal primordial gravity waves (pGWs) with left- and right-helicities [10, 11, 12, 13, 14, 15]. Thus, detecting either the helical asymmetry between two chiral-pGWs or their imprints on CMB polarization [16, 17, 18, 19] and LSS [20, 21, 22, 23, 24] could confirm the primordial tensorial-PV. Such helical asymmetry imprints would be more detectable in 3-dimensional probes [21] since the helicity information is largely lost in 2 dimensions, for example, leaving the detection in the polarized CMB [18, 25, 26, 27] more challenging [c.f. see 28, 29, 30, hinting at the non-vanishing EB-correlation of non-primordial origin, e.g., cosmic birefringence].
@@
Another interesting, but largely unexplored, class of parity violation is vectorial type, possibly generating helically asymmetric vector fossils [20] that may leave imprints in initial vector quantities. Recently, a 3-dimensional estimator “initial spin” (rigorously defined in Eq. 9) – a vector representing the spinning of a mass element torqued by the initial tidal field [31, 32, 33] – was proposed [34, 35] and examined [36] to search for vectorial helical asymmetry.
This proposal relies on two key properties of the initial spin field — 1) a vector field that only consists of left- and right-helical modes, and 2) its strong alignments with “galaxy spin” and “halo spin”, or the angular momentum direction of a galaxy [35, 37] and a dark matter halo [38, 39, 40, 34] at late time. The latter property suggests that the characteristics of the initial spin field will remain in the galaxy spin field. Therefore, detecting helical asymmetry in a galaxy spin field would indicate that the initial spin field is helically asymmetric, evidencing vectorial-PV in the early Universe.
@@
On the observational side, there has not yet been compelling evidence that the galaxy spin field is helically asymmetric. It has only been shown that there are mildly uneven correlations to the left- and right-helical components of the reconstructed initial spin field [36]. However, it is not straightforward to conclude that primordial PV is absent due to such non-detection. This is because galaxies’ nonlinear formation and evolution may have substantially diminished the primordial asymmetry, leaving only a weak or near-parity symmetric signal in galaxy spins. Thus, it is necessary to further investigate into how effectively the primordial helical asymmetry in the initial spin is carried onto the late-time galaxy spin.
In this Letter, we investigate the detectability of PV in a primordial vector fossil using galaxy spins. Specifically, we examine the PV in a “helical” (i.e., rotational) displacement vector field. The helical displacement vector field is envisioned to be produced in the very early universe, e.g., during inflation, and remains as fossils [20]. Note that the initial vector mode perturbations decay over time but can leave vector fossils [^1] We confirmed that the helical displacement remains a fossil in the linear regime, as its effect on the spin field at a fixed time is consistent across simulations initialized at different redshifts, all with the same fixed amplitude of the helical displacement., in analogy to tensor perturbations imprinting tensor fossils [41, 42, 21].
@@
With the help of numerical N-body simulations, we analyze halo spins as a reliable proxy for galaxy spins, leveraging the tight correlation between those two spins [43, 44, 40, 37]. In the spirit of effective field theory, without prior assumptions on microscopic physics, we develop a novel way of generating the initial condition with a helically asymmetric initial spin field. We show that the parity-broken displacement field efficiently produces strong primordial helical asymmetry without affecting the initial matter power spectrum. Our analysis of halos evolved from these initial conditions reveals that this asymmetry survives significantly in the final halo spin field.
@@
Simulations and initial conditions.— We construct 500 pairs of N-body simulations – ones with parity symmetric initial condition and the others with parity asymmetric initial condition. The simulations adopt the modified version of the code CUBE [45] to evolve $256^{3}$ dark matter particles within a periodic cubic box of size $L=100~h^{-1}{\rm Mpc}$. For numerical convenience, the initial redshift is set $z_{\rm init}=30$. We generate the initial density fluctuation of the standard parity symmetric universe using Zel’dovich approximation [46] where the displacement vector field, $\bm{\Psi}_{\rm ZA}$, is curl-free (i.e., irrotational) and its Fourier space representation follows
$\bm{\Psi}_{\rm ZA}=-i\bm{k}\frac{\Phi(\bm{k})}{4\pi Ga^{2}\bar{\rho}}\ ,$ (1)
@@
with the gravitational potential $\Phi$, scale factor $a$, gravitational constant $G$, and mean cosmic density $\bar{\rho}$. On the other hand, the parity asymmetric initial condition is constructed using the modified displacement field, $\bm{\Psi}_{\rm tot}$, which includes an extra helical displacement, $\bm{\Psi}_{\rm hel}$,
$\bm{\Psi}_{\rm tot}=\bm{\Psi}_{\rm ZA}+\bm{\Psi}_{\rm hel}.$ (2)
@@
$\bm{\Psi}_{\rm hel}$ $=\bm{\Psi}_{\mathcal{R}}+\bm{\Psi}_{\mathcal{L}}=\sum_{\mathcal{H}=\mathcal{R},\mathcal{L}}\Phi_{\mathcal{H}}(\bm{k})\hat{\bm{e}}_{\mathcal{H}}(\bm{k}).$ (3)
The $\mathcal{R}$- and $\mathcal{L}$-helical displacement fields can be constructed via $\bm{\Psi}_{\mathcal{R}/\mathcal{L}}=\Phi_{\mathcal{R/L}}\hat{\bm{e}}_{\mathcal{R/L}}$, with a pseudo-scalar field, $\Phi_{\mathcal{R}/\mathcal{L}}\equiv\hat{\bm{e}}_{\mathcal{R/L}}\bm{\Psi}_{\rm tot}$, and an eigenvector,
$\hat{\bm{e}}_{\mathcal{R}/\mathcal{L}}(\bm{k})=\frac{1}{\sqrt{2}k}\frac{1}{\sqrt{k_{y}^{2}+k_{z}^{2}}}\begin{pmatrix}k_{y}^{2}+k_{z}^{2}\\ -k_{x}k_{y}\pm ikk_{z}\\ -k_{x}k_{z}\mp ikk_{y}\end{pmatrix}\$, (4)
with $k=|\bm{k}|=\sqrt{k_{x}^{2}+k_{y}^{2}+k_{z}^{2}}$. The eigenvectors are obtained by solving the eigenvalue problem for the curl operator in Fourier space. The parity symmetry in the primordial vector fossil is broken if the modified displacement field includes two helical components unequally. To simplify our analysis, we assume $\bm{\Psi}_{\rm hel}$ to be purely right-handed, e.g., $\bm{\Psi}_{\rm tot}=\bm{\Psi}_{\rm ZA}+\bm{\Psi}_{\mathcal{R}}$. Note that both parity symmetric and asymmetric simulations should have identical initial density perturbation in Lagrangian space. This is because the initial density perturbation is solely generated by the same $\bm{\Psi}_{\rm ZA}$ [47], but not by $\bm{\Psi}_{\rm hel}$, as the latter is divergence-free. The helical displacement only re-maps the density perturbation at one (Eulerian) position to a different position [^2] The initial particle velocities in both simulations are purely set by the Zel’dovich approximation since any helical component eventually decays with expansion. and plays a pivotal role in breaking parity symmetry in the initial spin vector field, as we will show later.
We now describe the generation of $\Phi$ and $\Phi_{\mathcal{R}}$. In Fourier space, the density fluctuation, $\delta(\bm{k})$, is first constructed to follow the input matter power spectrum, $P(k)$,
$(2\pi)^{3}\delta_{D}(\bm{k}-\bm{k}^{\prime})P(k)=\langle\delta(\bm{k})\delta^{*}(\bm{k}^{\prime})\rangle,$ (5)
where its normalization amplitude is defined
$\sigma^{2}\equiv\frac{1}{2\pi^{2}}\int k^{2}P(k)dk,$ (6)
and set to $\sigma(z_{\rm init})=0.20$. Note that this corresponds to familiar $\sigma_{8}(z=0)=0.81$, consistent with CMB measurements [48]. We then obtain $\Phi(\bm{k})$ from $\delta(\bm{k})$ using the Poisson equation $k^{2}\Phi=4\pi Ga^{2}\bar{\rho}\delta$. Similarly, the pseudo-scalar field $\Phi_{\mathcal{R}}$ is obtained to satisfy
$\left\langle\left|\frac{k^{2}\Phi_{\mathcal{R}}}{4\pi Ga^{2}\bar{\rho}}\right|^{2}\right\rangle=P_{\mathcal{R}}(k),$ (7)
where we consider $P_{\mathcal{R}}(k)$ with high-$k$ suppression on $P(k)$ as
@@
We set $R_{\rm h}=0.6~h^{-1}{\rm Mpc}$ in correspondence to the Gaussian radius of a Lagrangian protohalo of $10^{12}~h^{-1}{\rm M}_{\odot}$. This is to minimize numerical artifacts generated by the helical displacement on scales smaller than the halos we investigate. The normalization is set to $\sigma_{\mathcal{R}}(z_{\rm init})=1.45$ to produce large enough parity asymmetry to be detected while maintaining the initial matter power spectrum unchanged. Qualitatively, $\sigma_{\mathcal{R}}$ determines the degree of primordial parity asymmetry, in analogy to non-gaussianity parameter, $f_{\rm NL}$, describing the departure from Gaussianity [49]. A universe without vectorial-PV corresponds to the case with $\sigma_{\mathcal{R}}=0$.
Independent random number sets are used for generating $\Phi$ and $\Phi_{\mathcal{R}}$ to make them uncorrelated. We identify halos using the Friends-of-Friends algorithm and utilize only halos more massive than $10^{12}~h^{-1}{\rm M}_{\odot}$ at $z=0$. The smallest mass halo contains at least $230$ particles, allowing for a reliable angular momentum measurement [50]. The total number of such halos in the entire 500 simulations of each kind is $N_{\rm halo}\simeq 10^{6}$.
Figure 1: (Eulerian) Initial matter density fields at $z_{\rm init}=30$ in the parity symmetric (left) and asymmetric (right) simulations sliced in a $0.4~h^{-1}{\rm Mpc}$-thickness slab. With the helical displacement (right), matter distribution looks rotationally distorted from that without it (left). Density fields are calculated using the cloud-in-cell (CIC) assignment scheme and smoothed with a Gaussian kernel over $R_{\rm h}=0.6h^{-1}{\rm Mpc}$. Note that they look identical in Lagrangian space.
Initial density field.— We begin by visually comparing the initial density fields of the parity symmetric and asymmetric simulations in Fig. 1. It can be seen that the helical displacement field $\Psi_{\mathcal{R}}$ contributes to distorting the standard density field similarly to the gravitationally lensed map of the CMB temperature and polarization [51, 52, 53]. For example, the density structures in the parity asymmetric realization tend to be stretched and/or warped from their counterparts at similar locations in the standard parity symmetric realization. However, unlike CMB maps exhibiting concentric distortion around over/under-densities, the density field deformation due to the helical displacement occurs incoherently.
Initial and halo spin fields.— We now examine the parity symmetry of the initial conditions. Specifically, we measure the asymmetry between the $\mathcal{R}$- and $\mathcal{L}$-helical components of the initial spin vector field. The initial spin field, $\bm{J}_{\rm init}$, is calculated on $256^{3}$ grids as a unit vector field constructed from the Hessians of the smoothed initial density, $\rho_{\rm init}^{s}$, and potential field, $\phi_{\rm init}^{s}$, [38, 54, 34, 35, 36],
@@
with the Levi-Civita symbol $\epsilon_{\alpha\beta\gamma}$. The optimal smoothing scale of the Gaussian kernel is set to $R_{\rm s}=2R_{\rm h}$, suppressing fluctuations on scales smaller than halos [35, 36, 34]. Note that the initial spin vector field is divergence-free, or purely helical, by construction. On the other hand, we construct the halo spin field, $\bm{J}_{\rm halo}$, by assigning halo spin vectors, $\bm{j}_{\rm halo}$, to the grids nearest to their respective protohalo centers, $\bm{x}_{\rm ph}$. This is expressed as
$\bm{J}_{\rm halo}(\bm{x})=\begin{cases}\bm{j}^{i}_{\rm halo}&\text{if $\bm{x}=\bm{x}_{\rm ph}^{i}$ for $i^{th}$ halo}\\ \bm{0}&\text{otherwise}.\end{cases}$ (10)
A halo spin vector represents the unit angular momentum of a halo at $z=0$ and is defined as,
$\bm{j}_{\rm halo}\equiv\frac{\sum_{j}\bm{r}_{p}^{j}\times\bm{v}_{p}^{j}}{|\sum_{j}\bm{r}_{p}^{j}\times\bm{v}_{p}^{j}|},$ (11)
where the numerator is the halo angular momentum, calculated as the sum of the angular momenta of its member particles. Here, $\bm{r}_{p}$ and $\bm{v}_{p}$ are particle position and velocity vectors relative to the halo center. Similarly, we construct the subsampled initial spin field using only the initial spin vectors that are assigned to the grids nearest to the protohalo centers. Thus, both the halo spin and subsampled initial spin fields have an equal number of nonzero spin vectors at identical positions.
Measuring helical asymmetry.— We examine the parity symmetry of the initial, subsampled initial, and halo spin fields by measuring the differences between their $\mathcal{R}$- and $\mathcal{L}$-handed components. We decompose each of the constructed spin fields, $\bm{J}$, into its $\mathcal{R}$- and $\mathcal{L}$-helical fields,
$\bm{J}_{\mathcal{R/L}}=(J_{\mathcal{R/L}})_{\alpha}=\sum_{\beta}\mathbb{P}^{\mathcal{R/L}}_{\alpha\beta}({\bm{k}})J_{\beta}({\bm{k}}),$ (12)
@@
The helical asymmetry, $\chi(k)$, of $\bm{J}$ is quantified as
@@
where $P_{{\mathcal{RR}/\mathcal{LL}}}(k)$ represents the power spectrum of the $\mathcal{R}$$\mathcal{L}$-helical spin field. It is defined as
@@
A spin vector field with perfect helical symmetry satisfies $\chi=0$, whereas $|\chi|$ equals unity when the spin field is maximally asymmetric, or purely $\mathcal{R}$$\mathcal{L}$-helical. Note that the cross-power spectrum of $\mathcal{R}$- and $\mathcal{L}$-helical fields vanishes because they are orthogonal.
Figure 2: Average helical asymmetry spectrum (upper) of initial spin fields and the significance of the departure (lower) from the perfect symmetry in the parity symmetric and asymmetric simulations. The vertical line marks a wavenumber corresponding to twice the Gaussian radius of a $10^{12}~h^{-1}{\rm M}_{\odot}$ protohalo, $k=1/(2R_{\rm h})$, whereas the horizontal line indicates $\chi/\sigma=50$. Shaded regions represent the standard error of the mean calculated from the 500 realizations, or the total volume of $0.5(h^{-1}{\rm Gpc})^{3}$.
In Fig. 2, we show the average helical asymmetry in the full initial spin fields for the parity symmetric/asymmetric realizations. We focus only on scales larger than $k=1/(2R_{\rm h})$, below which the helical asymmetry feature in halo spins is likely to be convolved with the halo exclusion effect [55]. In the parity asymmetric realizations, the initial spin field exhibits scale-dependent helical asymmetry that deviates from $\chi=0$, yielding about $70\sigma$ departure at maximum at $k\approx 0.3h{\rm Mpc}^{-1}$. Overall, the initial spin field of the parity asymmetric realization is preferentially $\mathcal{R}$-handed, e.g., $\chi>0$. In particular, the helical asymmetry can be as large as $\chi_{\rm peak}\approx 0.18$ indicating that the $\mathcal{R}$-helical mode is stronger by $44\%$ than the $\mathcal{L}$-helical mode. Such an asymmetry peak becomes higher and shifts toward a larger wavenumber if the initial spin field is constructed adopting a larger smoothing scale. On the other hand, regardless of the smoothing scale, we observe a nearly perfect helical symmetry (i.e., $\chi\approx 0$) on all scales investigated in the parity symmetric simulations. It is worth noting that the net helicity of the displacement field determines the relative dominance between the two helical modes of the initial spin fields.
@@
We now compare the late-time halo spin field with the subsampled initial spin field in Fig. 3. Unlike the full initial spin field, both spin fields contain shot-noise contamination because a finite number of spin vectors are used for their construction. To compensate for the shot-noise effect, we rescale by multiplying the same rescaling factor to their helical asymmetry spectra. Consequently, the subsampled and full initial spin fields for the parity asymmetric case have identical peak height, e.g., $\chi_{\rm peak}\simeq 0.18$. In the parity symmetric cases, the subsampled initial field shows larger fluctuations around $\chi=0$ than in the full initial spin field. Nonetheless, it is still consistent with the helical symmetry within $1.4\sigma$ significance. Similarly, we do not detect any significant deviation from the perfect symmetry in the halo spin field. The departure at maximum is only about $0.7\sigma$-level across the scales investigated. This indicates that the halo spin field remains parity symmetric if the initial spin field were parity symmetric.